### v1

@@ Discussion
Figures 3 d and e and panel 4 d show an clear agreement with each other. In the four cases (IAS-JT, IS-JT, M, and ), it is possible to see that the change in TB begins when the system is cycled near TC. As TT moves away from TC, the difference between the configurations before and after each cycle begins to diminish. Regarding the intensity of the JT modes and , a distinct dependence on TMI is evident, which is not prominently observed in the behavior of magnetization. For TT TMI, the difference between the initial and final states of the cycles tends towards zero, indicating the onset of the stable regime, as evidenced in previous figures, specifically panels 2 c and 4 c. The symmetry exhibited by the behavior of M and around TCOO, panel 4 d, underscores the validity of the prior interpretation of TCOO as the system’s reset temperature.
@@
Hence, Raman spectroscopy indicates both irreversible and reversible shifts in the FMM cluster distribution within the AFMI-COO matrix. An increase in resistivity correlates with reduced magnetization and heightened JT distortion modes, more pronounced in the insulating phase due to JT-polaronic transport and charge localization [Billinge2002, Kim]. This suggests a reduction, stabilization, and eventual return of the clusters to initial values at TB as a function of TT. More precisely, within the widely known percolation scenario for this compound [Uehara], thermal cycling changes the dimensionality of the percolation paths formed at TB according to the material’s thermal history. Since the change in the resistivity of the system is represented by a maximum increase (at TC TT TCOO) of about three times, it cannot be inferred that the pre-formed percolation paths are completely destroyed during the initial cooling of the sample to TB, prior to the execution of thermal cycles. The formation of these, when cooling from 300 K, can be seen in panel 4 b, with the 103 resistivity drop after the TIM transition at around 120 K.
@@
In figure 5 a semi-quantitative and illustrative interpretation of the DPS reconfiguration scenario, as a function of the target temperature, proposed by all the presented data, is constructed. As the temperature is varied cyclically, the relative size of the FMM regions, percolated among themselves, begins to decrease at TB when TT exceeds TC. In this process, it is also possible that a partial disruption of some percolation paths occurs. For values of TT between TC and TCOO, the behavior stabilizes, so that the decrease in the FMM regions ceases. This is exemplified in figure 5 c for TT = T and T. For TT TCOO the behavior is reversed, T, and the initial regime at TB is recovered, TT T.
@@
Our results suggest the occurrence of a non-volatile resistive-switching (RS) in the sample as a consequence of cumulative thermal cycling around the TC TT TMI region. In contrast to conventional RS studies, where such effects are typically induced by current and/or voltage, the observed phenomenon here is entirely thermally driven, utilizing the material’s inherent transitions for both irreversible and reversible modulation of resistivity. Moreover, for the first time, it is demonstrated how the resistive switching (RS) phenomenon can be associated with the local structure of the sample. More specifically, this switching is found to involve alterations in the degree of structural order and the relative abundance of the antisymmetric (AS) and symmetric Jahn-Teller (S-JT) distorted regions. The ability to induce these effects through localized heating via a light source, while concurrently analyzing the resulting changes with Raman spectroscopy, holds substantial significance for the advancement of resistive devices and neuromorphic computing. This is attributed to the sample’s capacity to store information structurally, magnetically, and electrically. Furthermore, it introduces an innovative method for controlling the phase separation (PS) regime and inducing resistive switching (RS) in complex oxides. Additional investigations parallel to those presented in this study, particularly focusing on manganites or other strongly correlated oxides known to exhibit resistive-switching (RS) responses, are highly desirable. This is especially true for studies where light serves as the primary mechanism of action.
@@ Conclusions
In conclusion, the study demonstrates the potential of laser-induced thermal cycling to induce non-volatile structure/resistive-switching in complex oxides, leveraging the intrinsic phase transitions of the material. It was observed for the that the resistive-switching effect is thermally driven and correlates with changes in structural distortion order. This work introduces a novel method for controlling phase separation and resistive properties through localized laser heating and characterizing the previously laser-heated regions by Raman spectroscopy. Such findings have significant implications for the development of advanced resistive devices and neuromorphic systems, showcasing the material’s ability to store and manage information structurally, magnetically, and electrically. Future investigations in similar oxide materials could further elucidate the role of localized heating and structural modulation in resistive-switching phenomena.
Methods
Synthesis and Synchrotron X-Ray Diffraction.
Polycrystalline La5/8-yPryCa3/8MnO3 () samples were prepared using a standard solid-state reaction method, as described by [li2008synthesis, Diego24]. Its orthorhombic Pnma (a c b) single phase character at room temperature was confirmed by synchrotron x-ray powder diffraction performed in the PAINEIRA beamline in CNPEM-SIRIUS [PAINEIRA, Collado2003]. The Rietveld refinement performed with the GSAS-II software [GSAS] and the refined crystallographic parameters can be found in figure S7 and table S2 in Supplementary Materials S.
Raman Spectroscopy
Raman scattering experiments were performed in a quasi-backscattering geometry using a 532 nm laser line with a spot size at the sample of 50 m. A Jobin-Yvon T64000 spectrometer with a single 600 mm-1 grating and a set of Bragg notch filters was used. A LN2-cooled multichannel charge-coupled device (Symphony II, Horiba) collected the scattered data. The sample was glued with silver paint on the cold finger of a closed-cycle He cryostat with a minimal temperature of K. The temperature was controlled by a LakeShore Model 340 temperature controller. The Raman’s setup is depicted in figure 1 a. Spectra as a function of the cold-finger nominal temperature were collected while heating to investigate the thermal evolution of the sample’s vibrational modes and to serve as a control parameter for the measurements performed as a function of laser power. This protocol will be referred to as ”standard procedure/heating”. In the measurements as a function of laser power, the sample was first cooled and maintained at the cryostat minimal temperature. Then, two different protocols were employed: 1) a continuous increase in power, with spectra collection using 10 mW to 150 mW (with 10 mW steps), and 2) cycling the power, where spectra were collected at 10 mW (base power) before and after it was increased to target powers ranging from 20 mW to 150 mW. The latter method of temperature variation (by the laser power), illustrated in Fig. 1 b, allows for almost instantaneous heating and subsequent freezing of the analyzed region, and will be referred as ”cyclic procedure/heating”. The comparison between the spectra collected in protocol 1) and those collected in the standard procedure yielded the temperature of the analyzed spot as a function of power, as detailed in the Supplementary Materials S. The thermal cycles were conducted between TB = 38(4) K and TT = 59(4), 81(4), …, 317 (7) K; see table S1 in the Supplementary Materials S. Recently broken, fresh, surfaces of the sample were measured. A 5-minute wait before spectral collection ensured thermal equilibrium on each measurement. In both experiments, the laser spot probed the same position on the sample to enhance control over the localized heating effects induced by the laser, and the collection time was proportionally decreased as the laser power increased (based on the 5 min collection time for 10 mW). Different spots and pieces of the sample were measured with the same protocol.
Magnetic and Electric properties
Magnetic and electric characterizations were performed using a PPMS Model 6000 (Quantum Design). Magnetization as a function of temperature, M-T, was measured in the ACMS mode with an AC-field amplitude of 10 Oe and an oscillating frequency of 11 Hz. Extractions were performed between 300 K and 10 K, cooling and warming at a rate of 2 K/min with no applied DC-field. The sample resistance was measured using a standard two-probe technique. Gold wires were glued with silver paint directly onto an LPCMO pellet, with spacing between analyzer probes of 0.7 mm. A fixed and constant applied current of 0.1 A was used. Resistance as a function of temperature, R-T, was measured between 300 K and 10 K, cooling and heating at a rate of 2 K/min, respectively. Magnetization and resistance as a function of the cyclic procedure, depicted in figure 1 a, between a base TB = 30 K and target temperatures, TT, 50, 60, 70, … 250 K, were measured after cooling to 30 K. Warming and cooling M-T and R-T curves between TB and TT were collected with a cooling/warming rate of 5 K/min.
Acknowledgments
@@
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001 and by the São Paulo Research Foundation (FAPESP), grant No. 2025/16064-5, 2024/00998-6 and 2023/17024-1. We also acknowledge the support of the INCT project Advanced Quantum Materials, involving CNPq (Proc. 408766/2024-7), FAPESP, and CAPES. This research used the facilities of the Brazilian Synchrotron Light Laboratory (LNLS), part of the Brazilian Center for Research in Energy and Materials (CNPEM), a private non-profit organization under the supervision of the Brazilian Ministry of Science, Technology, and Innovations (MCTI). The PAINEIRA beamline staff, especially Dra. Flávia Regina Estrada, are acknowledged for their assistance during the experiments 20250744. J.G. Ramirez and D. Carranza-Célis acknowledge support from Facultad de Ciencias and Vicerrectoría de investigaciones Universidad de los Andes, Convocatoria Facultad de Ciencias, Project No. INV-2023-178-2978. G. B. Gomide acknowledges the support of FAPESP under grant No. 2024/03819-5. M. Knobel and J.G. Ramirez acknowledge the support of FAPESP under grant No. 2022/16626-5.
S Supplementary Materials.
S Local temperature as a function of laser power
Raman spectra as a function of temperature and laser power were collected for the sample = 0.35. The local sample temperature as a function of laser power was determined by directly comparing spectra at a fixed power (10 mW) with variable temperatures and at a fixed temperature (16 K) with variable powers. These will be referred to as spectra as function of temperature and power, respectively. This comparison was based primarily on three factors: the similarity between the low-frequency regions, which is directly related to the thermal Bose-Einstein factor (and therefore to temperature); and the width and relative intensity of the three main spectral structures, centered at 250 , 480 , and 614 . After an initial estimation, the fit parameters of the three structures from both sets of data were compared to verify their agreement. It is assumed that the increase in local temperature due to the increase in power should be linear or approximately linear.
Figures S1 (a) and (b) present intensity maps corresponding to the spectra as functions of temperature and power, respectively. Within these maps, the principal inflection points in the intensities of the structures at 480 and 614 are delineated by dashed horizontal lines. In Figure S1 (a), it is apparent that there is an increase in the intensity of these structures at approximately 75 K. The maximum intensity is reached around 100 K. Then, after 125 K, the intensities of both structures begin to decrease. At 175 K there is an almost complete loss of intensity by the structure 614 and a large broadening, together with the loss of intensity, of the structure at 480 . In Figure S1 (b), it is possible to note that similar intensity inflections in the same structures of interest are observed at 40, 50, 80, and 100 mW, respectively. Therefore, it is possible to observe that there is a correspondence between the intensities of the mentioned regions for both heating protocols. It is important to note that for the data as a function of power, the intensity variation was accompanied by an increase in the background signal, which was not the same for all measurements. This, as well as possible increases in the intensity of elastic scattering, was also considered.
In figures S2 (a) and (b), some of the Raman spectra, vertically shifted, collected as a function of the heating produced by the cold finger heater and the increase in laser power, respectively, are shown. The spectra aligned horizontally in the two figures are those that present the greatest similarity to each other, considering the factors mentioned previously. The structures around 250, 485, and 614 cm-1 are identified in the figures by vertical dashed lines, while the low-frequency region ( - 300 cm-1 to 200 cm-1) is delimited by vertical dotted lines.
| Power (mW) | Cryo. Temp. (K) | Temp. Fit. (K) | Adj. Cryo Temp. (K) | Adj. Cryo Temp. Fit (K) |
| 10 | 16 | 20.9 (2) | 37 | 38 (4) |
| 20 | - | 41.5 (4) | - | 59 (4) |
| 30 | - | 62.8 (6) | - | 81 (4) |
| 40 | 75 | 83.7 (8) | 96 | 102 (4) |
| 50 | 100 | 105 (9) | 121 | 124 (5) |
| 60 | 125 | 126 (1) | 146 | 145 (5) |
| 70 | 150 | 146 (1) | 171 | 167 (5) |
| 80 | 175 | 167 (2) | 196 | 188 (5) |
| 90 | - | 188 (2) | - | 209 (6) |
| 100 | 200 | 209 (2) | 221 | 231 (6) |
| 110 | 225 | 230 (2) | 246 | 252 (6) |
| 120 | 250 | 251 (2) | 271 | 274 (7) |
| 130 | 275 | 272 (3) | 296 | 295 (7) |
| 140 | 300 | 293 (3) | 321 | 317 (7) |
| 150 | - | 314 (3) | - | 338 (8) |
Hence, from the figure, it is feasible to establish the following equivalence between temperature and power: 40 mW corresponds to 75 K; 50 mW to 100 K; 60 mW to 125 K; 70 mW to 150 K; 80 mW to 175 K; 100 mW to 200 K; and 140 mW to 300 K. However, the groups of spectra do not align perfectly, introducing some degree of uncertainty in this determination, particularly as not all spectra exhibit identical backgrounds. These uncertainties are accounted for in the linear fit.
For temperatures and powers lower than 75 K and 40 mW, respectively, no differences(except for small variations in the background and in the intensity of elastic scattering) are observed. Consequently, it can be inferred that, within the power range of 10 mW to 40 mW , the heating effect spans from 16 K to 50 K. Spectroscopic analyzes conducted at 16 K, 30 K, and 50 K corroborate the absence of distinguishable differences in the obtained spectra at these temperatures.
The complete power-temperature designation is organized in table S1. First, based on the first two columns of table S1,an approximately linear increase in temperature is observed with increasing laser power, and the data can be appropriately fitted to a linear equation. The best fit was obtained with the equation T(P) = 2.09 (0.02) P 222If both fitting parameters were varied, the intercept becomes negative, more precisely -3 (5) K. Since this does not make physical sense (zero power implies negative temperatures), it was fixed to zero., as shown in figure S3 (a). Substituting the values of power, P, into this equation, one obtains the third column of table S1.
At first glance, one might assume that 10 mW does not generate any heating at the analyzed point. Nevertheless, it can be observed that under the approximation of a linear relationship, a power of P = 110 mW corresponds to a temperature of T = 232 K. Thus, T / P = 2.18 K/mW, which implies that 10 mW is equivalent to a heating of 21.8 K. Therefore, it is necessary to add this value to the temperatures provided by the temperature controller. Thus, the fourth column of table S1 is obtained. Fitting these points to a linear equation, the best fit is T(P) = 2.14 (0.04)P + 16 (4), shown in figure S3 (b). The values obtained from this last fit can be found in the last column of table S1. The errors in the third and fifth columns are based on the propagation of standard errors. The temperature values according to the linear fit, fifth column of table S1, will be adopted for the data as a function of power.
In figures S4 (a), (b) and (c) the centers of the modes at 250 cm-1, 486 cm-1 and 614 cm-1 are shown, corresponding to rotation (Rot.), anti-symmetric Jahn-Teller distortion (AS-JT) and symmetric Jahn-Teller distortion (S-JT) modes, respectively, for the spectra collected as a function of temperature and power. Good agreement in the behavior and energy values for the three peaks can be observed over the entire temperature range and within the fitting error. Furthermore, the centers of the three structures show similar dependencies on the sample transition temperatures, TFPS, at 93 K, and TOCO, at 205 K.
In figures S5 (a), (b) and (c), the FWHM of the modes at 250 cm-1, 486 cm-1, and 614 cm-1, respectively, is shown. It can be observed that the width behavior of both heating protocols presents good agreement, as well as their dependence on the sample transition temperatures. In figures S5 (d), (e) and (f) the areas of the same modes are shown, respectively. Again, it is possible to observe a great agreement in both behavior and values between both two procedures for the three structures.
The differences observed for T TOCO in the energy of the rotational mode at 250 cm-1 and for the S-JT distortion mode at 614 cm-1, as shown in figures S4 (c) and (f), and the discrepancies observed in the FWHM of the mode at 250 cm-1 for T TOCO and of the mode at 614 cm-1 for TFPS T TOCO, figure S5 (a), can be attributed to the differences in the nature of both warming protocols. However, a more detailed study is needed.
Figures S4 and S5 show that, in relation to the general behavior of the three main structures in the Raman spectrum of LPCMO ( = 0.35), the temperature calibration performed as a function of power is adequate.
S.1 Raman experiment
The whole frequency range was fitted using a fixed predefined 7th degree polynomial and a group of Lorentzian curves in a Python script. The elastic scattering and the Anti-Stokes contributions were considered part of the background. The fitting quality was judged based on the visual inspection of the model against the data. These can be found in figures S6 (a) - (i).
S.2 Additional Information
| Rietveld Refinement Parameters | |||||
| Rwp (%) | Rp (%) | ||||
| 10.49 | 8.1 | 1.26 | |||
| a (Å) | b (Å) | c (Å) | Volume (Å3) | ||
| 5.44318(1) | 7.67835(2) | 5.43884(1) | 227.3145(6) | ||
| Site | x | y | z | Frac. | Uiso |
| La | 0.02507(6) | 0.25000 | -0.0049(1) | 0.275 | 0.00565(6) |
| Pr | 0.02507(6) | 0.25000 | -0.0049(1) | 0.35 | 0.00565(6) |
| Ca | 0.02507(6) | 0.25000 | -0.0049(1) | 0.375 | 0.00565(6) |
| Mn | 0.0000 | 0.0000 | 0.5000 | 1.000 | 0.00273(9) |
| O1 | -0.0101(6) | 0.250000 | 0.4388(8) | 1.000 | 0.007(1) |
| O2 | 0.7152(7) | -0.0367(4) | 0.2847(7) | 1.000 | 0.0153(8) |

### v2

@@ Discussion
Figures 3 (e) and (f) and figures 4 (c) and (d) show an clear agreement with each other. In the four cases (IAS-JT, IS-JT, M, and ), it is possible to see that the change in TB begins when the system is cycled near TC. As TT moves away from TC, the difference between the configurations before and after each cycle begins to diminish. Regarding the intensity of the JT modes and , a distinct dependence on TMI is evident, which is not prominently observed in the behavior of magnetization. For TT TMI, the difference between the initial and final states of the cycles tends towards zero, indicating the onset of the stable regime, as evidenced in previous figures, specifically 3 (e) and (f) and 5 (c) and (d). The symmetry exhibited by the behavior of M and around TCOO, figures 5 (a) and (b), underscores the validity of the prior interpretation of TCOO as the system’s reset temperature.
@@
Hence, Raman spectroscopy indicates both irreversible and reversible shifts in the FMM cluster distribution within the AFMI-COO matrix. An increase in resistivity correlates with reduced magnetization and heightened JT distortion modes, more pronounced in the insulating phase due to JT-polaronic transport and charge localization [6, 25]. This suggests a reduction, stabilization, and eventual return of the clusters to initial values at TB as a function of TT. More precisely, within the widely known percolation scenario for this compound [52], thermal cycling changes the dimensionality of the percolation paths formed at TB according to the material’s thermal history. Since the change in the resistivity of the system is represented by a maximum increase (at TC TT TCOO) of about three times, it cannot be inferred that the pre-formed percolation paths are completely destroyed during the initial cooling of the sample to TB, prior to the execution of thermal cycles. The formation of these, when cooling from 300 K, can be seen in figure 4 (c), with the 103 resistivity drop after the TIM transition at around 120 K.
@@
In figure 6 a semi-quantitative and illustrative interpretation of the DPS reconfiguration scenario, as a function of the target temperature, proposed by all the presented data, is constructed. As the temperature is varied cyclically, the relative size of the FMM regions, percolated among themselves, begins to decrease at TB when TT exceeds TC. In this process, it is also possible that a partial disruption of some percolation paths occurs. For values of TT between TC and TCOO, the behavior stabilizes, so that the decrease in the FMM regions ceases. This is exemplified in figure 6 for TT = T and T. For TT TCOO the behavior is reversed, T, and the initial regime at TB is recovered, TT T.
@@
Our results suggest the occurrence of a non-volatile resistive-switching (RS) in the sample as a consequence of cumulative thermal cycling around the TC TT TMI region. In contrast to conventional RS studies, where such effects are typically induced by current and/or voltage, the observed phenomenon here is entirely thermally driven, utilizing the material’s inherent transitions for both irreversible and reversible modulation of resistivity. Moreover, it is demonstrated how the resistive switching (RS) phenomenon can be associated with the local structure of the sample. More specifically, this switching is found to involve alterations in the degree of structural order and the relative abundance of the antisymmetric (AS) and symmetric Jahn-Teller (S-JT) distorted regions. The ability to induce these effects through localized heating via a light source, while concurrently analyzing the resulting changes with Raman spectroscopy, holds substantial significance for the advancement of resistive devices and neuromorphic computing. This is attributed to the sample’s capacity to store information structurally, magnetically, and electrically. Furthermore, it introduces an innovative method for controlling the phase separation (PS) regime and inducing resistive switching (RS) in complex oxides. Additional investigations parallel to those presented in this study, particularly focusing on manganites or other strongly correlated oxides known to exhibit resistive-switching (RS) responses, are highly desirable. This is especially true for studies where light serves as the primary mechanism of action.
@@ Conclusions
In conclusion, the study demonstrates the potential of laser-induced thermal cycling to induce non-volatile structure/resistive-switching in complex oxides, leveraging the intrinsic phase transitions of the material. It was observed that the resistive-switching effect is thermally driven and correlates with changes in structural distortion order. This work introduces a novel method for controlling phase separation and resistive properties through localized laser heating and characterizing the previously laser-heated regions by Raman spectroscopy. Such findings have significant implications for the development of advanced resistive devices and neuromorphic systems, showcasing the material’s ability to store and manage information structurally, magnetically, and electrically. Future investigations in similar oxide materials could further elucidate the role of localized heating and structural modulation in resistive-switching phenomena.
@@
Acknowledgments – This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001 and by the São Paulo Research Foundation (FAPESP), grant No. 2025/16064-5, 2024/00998-6 and 2023/17024-1. We also acknowledge the support of the INCT project Advanced Quantum Materials, involving CNPq (Proc. 408766/2024-7), FAPESP, and CAPES. This research used the facilities of the Brazilian Synchrotron Light Laboratory (LNLS), part of the Brazilian Center for Research in Energy and Materials (CNPEM), a private non-profit organization under the supervision of the Brazilian Ministry of Science, Technology, and Innovations (MCTI). The PAINEIRA beamline staff, especially Dra. Flávia Regina Estrada, are acknowledged for their assistance during the experiments 20250744. J.G. Ramirez and D. Carranza-Célis acknowledge support from Facultad de Ciencias and Vicerrectoría de investigaciones Universidad de los Andes, Convocatoria Facultad de Ciencias, Project No. INV-2023-178-2978. G. B. Gomide acknowledges the support of FAPESP under grant No. 2024/03819-5. M. Knobel and J.G. Ramirez acknowledge the support of FAPESP under grant No. 2022/16626-5.